# Learning the Unlearned: Mitigating Feature Suppression in Contrastive Learning

### v2

@@ Abstract
Self-Supervised Contrastive Learning has proven effective in deriving high-quality representations from unlabeled data. However, a major challenge that hinders both unimodal and multimodal contrastive learning is feature suppression, a phenomenon where the trained model captures only a limited portion of the information from the input data while overlooking other potentially valuable content. This issue often leads to indistinguishable representations for visually similar but semantically different inputs, adversely affecting downstream task performance, particularly those requiring rigorous semantic comprehension. To address this challenge, we propose a novel model-agnostic Multistage Contrastive Learning (MCL) framework. Unlike standard contrastive learning which inherently captures one single biased feature distribution, MCL progressively learns previously unlearned features through feature-aware negative sampling at each stage, where the negative samples of an anchor are exclusively selected from the cluster it was assigned to in preceding stages. Meanwhile, MCL preserves the previously well-learned features by cross-stage representation integration, integrating features across all stages to form final representations. Our comprehensive evaluation demonstrates MCL’s effectiveness and superiority across both unimodal and multimodal contrastive learning, spanning a range of model architectures from ResNet to Vision Transformers (ViT). Remarkably, in tasks where the original CLIP model has shown limitations, MCL dramatically enhances performance, with improvements up to threefold on specific attributes in the recently proposed MMVP benchmark.
@@ 1 Introduction
Self-supervised contrastive learning obtains high-quality representations by maximizing the similarity between an anchor and its associated positive samples, while concurrently increasing the separation among the dissimilar data samples in the embedding space [22]. Various contrastive learning models serve as fundamental pretrained backbones across different fields [33, 6, 18, 43, 12, 13]. However, recent studies [36, 40, 45] have shown that representations derived from standard contrastive learning often miss substantial portions of input information. This phenomenon is referred to as feature suppression. Such suppression can severely compromise the effectiveness of models in various downstream tasks, ranging from classification [36] to pattern recognition [41]. In addition, the feature suppression issues are also observed in multimodal contrastive learning [3, 41], such as CLIP [33], which is predominantly adopted in current multimodal large language models(MLLMs) [29, 49, 46] as the vision encoder. Feature suppression in CLIP significantly impedes the capability of MLLMs to differentiate between images with varying semantics, resulting in severe hallucination problems [41, 27, 28, 26] within these models.
@@ 2.1 Contrastive Learning
How to extract useful information from unlabeled data is an important question in machine learning [2]. Among all branches of methods, contrastive learning flourishes in recent years [22] and plays an important role in text-to-image generation [33, 35, 34] and multimodal large language models [29, 49, 46]. Contrastive learning aims to learn useful representations from unlabeled data by maximizing the agreement between different views of the data [14]. Recently, different variants of contrastive learning methods have been proposed [6, 18, 32, 4]. CPC [32] learns representations by predicting future samples in a sequence using an auto-regressive model with a contrastive loss. MoCo [18] is designed to overcome the limitations of batch size in contrastive learning by introducing a dynamic dictionary with a queue and a moving-averaged encoder. SimCLR [6] is a simple yet effective framework for contrastive learning of visual representations. SimCLR demonstrated that, with sufficiently large batch sizes, it is possible to learn powerful representations without needing specialized architectures or a memory bank. MoCo-v2 [8] builds upon the original MoCo by incorporating several improvements from SimCLR. Negative-free contrastive learning [9, 16] further simplifies contrastive learning by removing the requirement of explicit negative samples. Although contrastive learning achieves promising performance in many fields [42, 47, 39, 43], it cannot guarantee all semantically relevant features are learned when multiple features exist [7, 44, 45, 3, 41].
@@ 2.2 Feature Suppression in Contrastive Learning
The feature suppression phenomenon has been first empirically observed by Chen et al. [7]. Subsequently, Robinson et al. [36] formally brings up this problem as feature suppression. A few prior works studied the cause of feature suppression. Robinson et al. [36] shows that simply minimizing the InfoNCE loss cannot avoid feature suppression. Both Robinson et al. [36] and Kukleva et al. [24] observed that the temperature parameter affects the trade-off of which features are learned and which features are suppressed. Xiao et al. [44] have discovered that certain augmentations used to generate positive samples might destroy the feature information, hence hindering the learning process of corresponding features. Assran et al. [1] point out that feature suppression might be caused by the hidden prior distribution bias in contrastive learning. Xue et al. [45] demonstrate that the simplicity bias of stochastic gradient descent is one of the factors. Not only in the above unimodal setting, Bleeker et al. [3] first studies this problem in the multimodal setting. Tong et al. [41] have observed that images with different semantics have unreasonably high similarities in CLIP [33] embedding space, which is also highly related to feature suppression in multimodal contrastive learning.
Few methods specifically address the challenge of feature suppression in contrastive learning. Robinson et al. [36] proposed a technique aimed at eliminating whichever features distinguish the positive sample from negative samples in the embedding space. Similarly, Tamkin et al. [40] applied a comparable strategy but targeted the input space. Both approaches are based on adversarial training, which does not ensure the preservation of previously well-learned features. In contrast, our approach diverges by sequentially learning new features stage by stage without compromising the integrity of already learned features. Bleeker et al. [3] mitigate the feature suppression problem by introducing an additional reconstruction loss, which is not feasible for large-scale settings such as CLIP due to the high computational cost. In contrast, our approach does not necessitate modifications to the original loss function or alterations to the base model, conserving computational resources and ensuring compatibility across different models. It is worth noting that, unlike the aforementioned methods that are confined to either unimodal or multimodal settings, our work stands as the first attempt to tackle feature suppression across both unimodal and multimodal contrastive learning.
@@ 3.1 Self-Supervised Contrastive Learning
(a) Linear evaluation results of SimCLR and MoCo-v2 trained on Trifeature and CIFAR-MNIST. The feature suppression problem in SimCLR is severe. In Trifeature, SimCLR significantly ignores the shape information, and in CIFAR-MNIST, it almost completely neglects the CIFAR information. With MoCo-v2, the issue of feature suppression is less pronounced but still exists, considering the two datasets are very simple.
@@
CIFAR-MNIST(CIFAR) 0.10 0.77
CIFAR-MNIST(MNIST) 0.99 0.98
@@ 3.2 Feature Suppression
[…]
@@ 5.1 Datasets
Trifeature. Trifeature [20] is an image dataset, where each image is $128\times 128$ with three independent features: color, texture and shape each taking 10 values. For each combination of color, texture, and shape, there are 100 samples, in which the position and rotation of the object are random. The three downstream tasks are to classify the color ($C=10$ classes), texture ($C=10$ classes), and shape ($C=10$ classes) of the image.
CIFAR-MNIST. CIFAR-MNIST consists of channel-wise concatenation of the CIFAR-10 [23] image and the MNIST [25] image, following Chen et al. [7]. Each image has four channels: three from the CIFAR-10 image and one from the MNIST image. As the images are randomly sampled from the two datasets in concatenation, the CIFAR-10 class and the MNIST class can be considered as two independent features. The two downstream tasks are to predict the CIFAR-10 class ($C=10$ classes) and the MNIST class ($C=10$ classes).
@@
STL-10. STL-10 [11] is an image recognition dataset suitable for developing feature learning algorithms with more than 100k images in total. Each image is $96\times 96$. The downstream task is to predict the image label ($C=10$ classes).
@@ 5.4 Mitigating Feature Suppression in the Unimodal Setting
Unlike CIFAR-MNIST and Trifeature, STL-10 does not have explicitly identified semantic features. We incorporate MCL with both SimCLR and MoCo-v2 [8], use ResNet-50 [19] as the encoder, train on STL-10 dataset for $3$ stages, $400$ epochs for each stage. The results are shown in Fig. 4. MCL boosts the MoCo-v2 performance on STL-10 by more than two percent. This further validates the effectiveness of our proposed MCL framework.
@@ 5.5 Mitigating Feature Suppression in the Multimodal Setting
Experimental Results. Table 3 presents the performance of the tuned CLIP models across different stages. For both the ResNet and ViT architectures, we observe improvements in later stages over earlier ones on tasks where initial performances were suboptimal. For example, ViT at Stage 1 outperforms its Stage 0 counterpart in PSF, ViT at Stage 2 exceeds the performance of both Stage 0 and Stage 1 in O&C, and ViT at Stage 3 surpasses the earlier stages in C&A. These improvements underscore the models’ evolving expertise in distinct features, enhancing their competence with specific attributes. Notice that each stage of the ViT model specializes in different attributes: ViT-STAGE0 in O&D, ViT-STAGE1 in PSF, S&C, and V&P, ViT-STAGE2 in Q&C, and ViT-STAGE3 in C&A and S&P.
@@ 5.6 Discussion
Refer to caption (a) Linear evaluation accuracy for each stage individually before cross-stage representation integration. Refer to caption (b) Linear evaluation accuracy for training $N$ stages after cross-stage representation integration.
@@
CIFAR-MNIST(CIFAR) 0.29 0.83 ($\uparrow$.54) 0.10 0.87 ($\uparrow$.77) 0.10 0.87 ($\uparrow$.77)
CIFAR-MNIST(CIFAR) 0.98 0.98 ($\uparrow$.00) 0.99 0.99 ($\uparrow$.00) 0.99 0.99 ($\uparrow$.00)
@@ 6 Conclusion
In this paperr, we investigated the critical feature suppression in contrastive learning. Specifically, we introduced the Multistage Contrastive Learning (MCL) framework, a novel, model-agnostic framework. MCL employs a cross-stage negative sampling strategy that effectively promotes the learning of previously unlearned information at each stage. Meanwhile, MCL efficiently preserves well-learned features and mitigates degradation observed in prior works. The effectiveness of our approach is demonstrated through comprehensive analyses with commonly used baseline models on various datasets and settings, highlighting MCL’s effectiveness and adaptability in both unimodal and multimodal contrastive learning.
@@ Acknowledgement
This project was funded by the National Research Foundation Singapore under the RIE2025 Industry Alignment Fund (I2101E0002—Cisco-NUS Accelerated Digital Economy Corporate Laboratory) and under AI Singapore Programme (Award Number: AISG-GC-2019-001-2A).

### v3

@@ Abstract
Self-Supervised Contrastive Learning has proven effective in deriving high-quality representations from unlabeled data. However, a major challenge that hinders both unimodal and multimodal contrastive learning is feature suppression, a phenomenon where the trained model captures only a limited portion of the information from the input data while overlooking other potentially valuable content. This issue often leads to indistinguishable representations for visually similar but semantically different inputs, adversely affecting downstream task performance, particularly those requiring rigorous semantic comprehension. To address this challenge, we propose a novel model-agnostic Multistage Contrastive Learning (MCL) framework. Unlike standard contrastive learning which inherently captures one single biased feature distribution, MCL progressively learns previously unlearned features through feature-aware negative sampling at each stage, where the negative samples of an anchor are exclusively selected from the cluster it was assigned to in preceding stages. Meanwhile, MCL preserves the previously well-learned features by cross-stage representation integration, integrating features across all stages to form final representations. Our comprehensive evaluation demonstrates MCL’s effectiveness and superiority across both unimodal and multimodal contrastive learning, spanning a range of model architectures from ResNet to Vision Transformers (ViT). Remarkably, in tasks where the original CLIP model has shown limitations, MCL dramatically enhances performance, with improvements up to threefold on specific attributes in the recently proposed MMVP benchmark. Codes are available at https://github.com/MajorDavidZhang/MCL.git.
Keywords:
Self-Supervised Learning Contrastive Learning Feature Suppression
@@ 1 Introduction
Self-Supervised contrastive learning obtains high-quality representations by maximizing the similarity between an anchor and its associated positive samples, while concurrently increasing the separation among the dissimilar data samples in the embedding space [23]. Various contrastive learning models serve as fundamental pretrained backbones across different fields [37, 6, 19, 50]. However, recent studies [40, 45, 53] have shown that representations derived from standard contrastive learning often miss substantial portions of input information. This phenomenon is referred to as feature suppression. Such suppression can severely compromise the effectiveness of models in various downstream tasks, ranging from classification [40] to pattern recognition [47]. In addition, the feature suppression issues are also observed in multimodal contrastive learning [3, 47], such as CLIP [37], which is predominantly adopted in current multimodal large language models(MLLMs) [32, 58, 54] as the vision encoder. Feature suppression in CLIP significantly impedes the capability of MLLMs to differentiate between images with varying semantics, resulting in severe hallucination problems [47, 30, 31, 29] within these models.
@@ 2.1 Contrastive Learning
How to extract useful information from unlabeled data is an important question in machine learning [2]. Among all branches of methods, contrastive learning flourishes in recent years [23] and plays an important role in text-to-image generation [37, 39, 38] and multimodal large language models [32, 58, 54]. Contrastive learning aims to learn useful representations from unlabeled data by maximizing the agreement between different views of the data [14]. Recently, different variants of contrastive learning methods have been proposed [6, 19, 36, 4]. CPC [36] learns representations by predicting future samples in a sequence using an auto-regressive model with a contrastive loss. MoCo [19] is designed to overcome the limitations of batch size in contrastive learning by introducing a dynamic dictionary with a queue and a moving-averaged encoder. SimCLR [6] is a simple yet effective framework for contrastive learning of visual representations. SimCLR demonstrated that, with sufficiently large batch sizes, it is possible to learn powerful representations without needing specialized architectures or a memory bank. MoCo-v2 [9] builds upon the original MoCo by incorporating several improvements from SimCLR. Negative-free contrastive learning [10, 17] further simplifies contrastive learning by removing the requirement of explicit negative samples. Although contrastive learning achieves promising performance in many fields [48, 55, 44, 50, 26, 27], it cannot guarantee all semantically relevant features are learned when multiple features exist [7, 52, 53, 3, 47].
@@ 2.2 Feature Suppression in Contrastive Learning
The feature suppression phenomenon has been first empirically observed by Chen et al. [7]. Subsequently, Robinson et al. [40] formally brings up this problem as feature suppression, and shows that simply minimizing the InfoNCE loss cannot avoid feature suppression. Both Robinson et al. [40] and Kukleva et al. [25] observed that the temperature parameter affects the trade-off of which features are learned and which features are suppressed. Xiao et al. [52] have discovered that certain augmentations used to generate positive samples might destroy the feature information, hence hindering the learning process of corresponding features. Assran et al. [1] point out that feature suppression might be caused by the hidden prior distribution bias in contrastive learning. Xue et al. [53] demonstrate that the simplicity bias of stochastic gradient descent is one of the factors. Not only in the above unimodal setting, Bleeker et al. [3] first studies this problem in the multimodal setting. Tong et al. [47] have observed that images with different semantics have unreasonably high similarities in CLIP [37] embedding space, which is also highly related to feature suppression in multimodal contrastive learning.
Few methods specifically address the challenge of feature suppression in contrastive learning. Robinson et al. [40] proposed a technique aimed at eliminating whichever features distinguish the positive sample from negative samples in the embedding space. Similarly, Tamkin et al. [45] applied a comparable strategy but targeted the input space. Both approaches are based on adversarial training, which does not ensure the preservation of previously well-learned features. In contrast, our approach diverges by sequentially learning new features stage by stage without compromising the integrity of already learned features. Bleeker et al. [3] mitigate the feature suppression problem by introducing an additional reconstruction loss, which is not feasible for large-scale settings such as CLIP due to the high computational cost. In contrast, our approach does not necessitate modifications to the original loss function or alterations to the base model, conserving computational resources and ensuring compatibility across different models. It is worth noting that, unlike the aforementioned methods and several other related approaches [35, 8, 12, 16, 43] that are confined to either unimodal or multimodal settings, our work stands as the first attempt to tackle feature suppression across both unimodal and multimodal contrastive learning.
@@ 3.1 Self-Supervised Contrastive Learning
(a) Linear evaluation results of SimCLR and MoCo-v2 trained on Trifeature and CIFAR-MNIST (C-M). The feature suppression problem in SimCLR is severe. In Trifeature, SimCLR significantly ignores the shape information, and in CIFAR-MNIST, it almost completely neglects the CIFAR information. With MoCo-v2, the issue of feature suppression is less pronounced but still exists, considering the two datasets are very simple.
@@
C-M(CIFAR) 0.10 0.77
C-M(MNIST) 0.99 0.98
@@ 3.2 Feature Suppression
In the unimodal setting, the ability of the encoder to capture specific feature information can be assessed through the linear evaluation accuracy of discriminating that feature. A high linear evaluation accuracy for a given feature suggests the encoder has successfully captured substantial information regarding that feature, and vice versa. We train ResNet-18 encoders using SimCLR [6] and MoCo-v2 [9] on two datasets (CIFAR-MNIST and Trifeature) [^1] More detailed dataset description and settings can be found in Sec. 5. following Robinson et al. [40] and Chen et al. [7], to demonstrate the feature suppression phenomenon. The linear evaluation results for each feature across both datasets are shown in Tab. 1(a). In CIFAR-MNIST, the linear evaluation accuracy for MNIST features is high but both SimCLR and MoCo-v2 show comparatively lower accuracy for CIFAR features. This discrepancy indicates MNIST features’ predominance, with CIFAR information being notably overlooked—SimCLR, in particular, demonstrates this trend more pronouncedly. In Trifeature, a similar pattern emerges: both contrastive learning methods achieve high linear evaluation accuracy for texture and color, yet falter when it comes to shape. This divergence suggests that the encoders while the encoders sufficiently capture texture and color, they neglect shape information. Consequently, as illustrated in Fig. 1(a), images with different shapes in Trifeature have high similarity in the SimCLR embedding space. This overlap significantly hampers the model’s capacity to discern shapes, adversely affecting performance on downstream tasks reliant on shape differentiation.
@@ 5.1 Datasets
Trifeature. Trifeature [21] is an image dataset, where each image has three independent features: color, texture and shape each taking 10 values. For each combination of the three features there are 100 samples, in which the position and rotation of the object are random. The three downstream tasks are to classify the color ($C=10$ classes), texture ($C=10$ classes), and shape ($C=10$ classes).
CIFAR-MNIST (C-M). CIFAR-MNIST consists of channel-wise concatenation of the CIFAR-10 [24] image and the MNIST [28] image, following Chen et al. [7]. Each image has four channels: three from the CIFAR-10 and one from the MNIST. As the images are randomly sampled from the two datasets in concatenation, the CIFAR-10 class and the MNIST class can be considered as two independent features. The two downstream tasks are to predict the CIFAR-10 class ($C=10$ classes) and the MNIST class ($C=10$ classes).
@@ 5.4 Mitigating Feature Suppression in the Unimodal Setting
Unlike CIFAR-MNIST and Trifeature, STL-10 [13] does not have explicitly identified semantic features. We incorporate MCL with both SimCLR and MoCo-v2 [9], use ResNet-50 [20] as the encoder, train on STL-10 dataset for $3$ stages, $400$ epochs for each stage. The results are shown in Fig. 4. MCL boosts the MoCo-v2 performance on STL-10 by more than two percent. This further validates the effectiveness of our proposed MCL framework.
@@ 5.5 Mitigating Feature Suppression in the Multimodal Setting
Experimental Results. Table 3 presents the performance of the tuned CLIP models across different stages. For both the ResNet and ViT architectures, we observe improvements in later stages over earlier ones on tasks where initial performances were suboptimal. For example, ViT at Stage 1 outperforms its Stage 0 counterpart in PSF, ViT at Stage 2 exceeds the performance of both Stage 0 and Stage 1 in Q&C, and ViT at Stage 3 surpasses the earlier stages in C&A. These improvements underscore the models’ evolving expertise in distinct features, enhancing their competence with specific attributes. Notice that each stage of the ViT model specializes in different attributes: ViT-STAGE0 in O&D, ViT-STAGE1 in PSF, S&C, and V&P, ViT-STAGE2 in Q&C, and ViT-STAGE3 in C&A and S&P.
@@ 5.6 Discussion
(a) Linear evaluation accuracy for each stage individually before cross-stage representation integration.
(b) Linear evaluation accuracy for training $N$ stages after cross-stage representation integration.
@@
C-M(CIFAR) 0.29 0.83 ($\uparrow$.54) 0.10 0.87 ($\uparrow$.77) 0.10 0.87 ($\uparrow$.77)
C-M(MNIST) 0.98 0.98 ($\uparrow$.00) 0.99 0.99 ($\uparrow$.00) 0.99 0.99 ($\uparrow$.00)
@@ 6 Conclusion
In this paper, we investigated the critical feature suppression in contrastive learning. Specifically, we introduced the Multistage Contrastive Learning (MCL) framework, a novel, model-agnostic framework. MCL employs a cross-stage negative sampling strategy that effectively promotes the learning of previously unlearned information at each stage. Meanwhile, MCL efficiently preserves well-learned features and mitigates degradation observed in prior works. The effectiveness of our approach is demonstrated through comprehensive analyses with commonly used baseline models on various datasets and settings, highlighting MCL’s effectiveness and adaptability in both unimodal and multimodal contrastive learning.
@@ Acknowledgement
This project was funded by the National Research Foundation Singapore under AI Singapore Programme (Award Number: AISG-GC-2019-001-2B and AISG2-TC-2022-004).
Appendix 0.A Appendix
In the appendix, we provide further details and additional experimental results that complement the main text. The contents include:
Section 0.A.1 . Implementation details of feature-aware negative sampling.
Section 0.A.2 . Learned features hinder the model from acquiring new features.
Section 0.A.3 . Train SimCLR using MCL on ImageNet.
Section 0.A.4 . Ensemble CLIP models trained using MCL with Model Soup.
Section 0.A.5 . The computational cost of MCL.
Section 0.A.6 . Ablation study for the number of trainable blocks when training ViT CLIP models using MCL.
Section 0.A.7 . Analysis of the clustering results across different stages of MCL.
0.A.1 Details of Feature-Aware Negative Sampling
The feature-aware negative sampling strategy necessitates that an anchor and its negative samples belong to the same cluster, as determined in the previous stages. To effectively organize training samples according to their pseudo labels obtained between stages, we utilize a custom batch sampler. The implementation details of the batch sampler are shown in Algorithm 1.
Algorithm 1 Feature-Aware Negative Sampling for MCL
1:Input: Dataset $D$ with pairs of data and pseudo labels $\{(x_{i},y_{i})\}_{i=1}^{M}$, number of unique pseudo labels $C=|\{y_{i}\}_{i=1}^{M}|$
2:Group $D$ into a two-dimensional array G based on pseudo labels $y$, such that $\texttt{G}[c]=\{(x_{i},y_{i})\in D|y_{i}=c\}$
3:for batch index $j\leftarrow 0$ to $\infty$ do
4:     Select group index $k\leftarrow j\mod C$
5:     if all samples in $\texttt{G}[k]$ are traversed then
6:         continue
7:     end if
8:     Create batch from untraversed samples in $\texttt{G}[k]$
9:     if all samples in G are traversed then
10:         break
11:     end if
12:end for
0.A.2 Learned Features Hinder the Learning Process
We conduct experiments to demonstrate that the learned features can impede the model’s ability to acquire new features. Specifically, at each stage, we initialize the model with parameters from the previous stage instead of training from scratch. We train SimCLR with ResNet18 as the encoder on the Trifeature dataset. We use the $\tau=0.5$ setting as described in Section 5.6 in the main paper. As shown in Tab. 5, this model makes almost no progress from Stage 1 to Stage 2, and its final performance (0.42/0.97/1.00) is markedly inferior to that of our proposed method (0.75/0.99/1.00). This can be attributed to the design of MCL, which does not require the retention of previously learned properties at each stage, thus allowing it to freely learn new features without constraints.
Table 5: Linear evaluation results on shape, texture, and color respectively using naive inheritance and MCL.
Stage0 Stage1 Stage2
Inheritance 0.44/0.92/1.00 0.40/0.97/1.00 0.42/0.97/1.00
MCL 0.44/0.92/1.00 0.72/0.97/1.00 0.75/0.99/1.00
0.A.3 Train SimCLR Using MCL on ImageNet
To further validate the effectiveness of MCL on the large unimodal dataset, we train SimCLR with ResNet34 as the encoder on ImageNet. We use a batch size of 8192 and train for 100 epochs at each stage for 3 stages. As shown in Tab. 6, the experiment results demonstrate a notable improvement.
Table 6: Linear evaluation accuracy of incorporating MCL with SimCLR on ImageNet.
Stage0 Stage0+1 Stage0+1+2
MCL 0.456 0.477 0.482
0.A.4 Ensemble ViT CLIP Models with Model Soup
We conduct experiments to compare MCL fine-tuned models with the standard OpenAI pretrained model, maintaining identical parameter sizes. Specifically, we utilize Model Soup [51], an effective and efficient ensemble approach that averages the weights of multiple fine-tuned models. We applied Model Soup to ensemble the fine-tuned ViT CLIP models across four stages, as described in Section 5.4 of the main text. As illustrated in Tab. 7, the MCL Model Soup achieves an average score of 29.6. Although this represents a slight decrease in performance compared to the simple concatenation method discussed in our main text, it still substantially surpasses the baseline average score of 20.0, with the same parameter size. We also test the MCL Model Soup on text-to-image and image-to-text retrieval tasks, following the settings in Zhang et al. [57]. As shown in Tab. 8, the MCL Model Soup achieves a notable performance improvement compared to the OpenAI baseline. The results underscore the robustness and effectiveness of the MCL framework.
Table 7: Comparative analysis of ViT CLIP models fine-tuned with MCL on the MMVP benchmark. “OpenAI” denotes the baseline performance using the OpenAI pretrained model. “MCL Concat” represents the performance after applying MCL and using concatenated representations from all stages. MCL Model Soup (denoted as “MCL MS”) illustrates the performance of the ensemble model created by Model Soup.
O&D PSF S&C Q&C P&R C&A S&P Texts V&P Average
OpenAI 26.7 13.3 26.7 6.7 6.7 40.0 26.7 13.3 20.0 20.0
MCL Concat 6.7 20.0 73.3 13.3 13.3 80.0 46.7 13.3 26.7 32.6
MCL MS 13.3 26.7 46.7 20.0 6.7 60.0 26.7 33.3 33.3 29.6
Table 8: Performance of MCL Model Soup. Tasks include text-to-image (T2I) retrieval and image-to-text (I2T) retrieval on 5k COCO validation set and 30k Flickr30k dataset. We use top-1, top-5, and top-10 Recall (R@1, R@5, R@10) as the evaluation metrics.
COCO Flickr30k
Image-to-Text Text-to-Image Image-to-Text Text-to-Image
R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
OpenAI 56.1 79.5 86.8 35.4 60.1 70.2 48.5 72.6 80.8 28.0 49.3 58.7
MCL MS 60.4 82.4 89.5 42.2 66.9 76.2 52.7 76.6 84.4 36.5 59.0 68.1
0.A.5 The Computational Cost of MCL
The computational cost of MCL is detailed in Tab. 9, using our largest model in unimodal (as described in Sec. 0.A.3) and multimodal settings for illustration. Notably, the fine-tuning cost of the CLIP model is approximately 3% of the OpenAI pre-training cost in TFLOPs, and the inference cost increases by only 75%.
Table 9: Computational complexity of MCL models. The training cost is evaluated by TFLOPs, Wall Clock Time (WCT), and TFLOPS. Inference cost is evaluated by relative complexity ratio (R) compared to backbone models (i.e. SimCLR ResNet34 and CLIP ViT). WCT is measured on 8 A100 GPUs.
TFLOPs WCT TFLOPS R
MCL SimCLR ResNet34 $1.3\times 10^{7}$ 30h $1.6\times 10^{2}$ 300%
MCL CLIP ViT $1.8\times 10^{8}$ 32h $1.6\times 10^{3}$ 175%
0.A.6 Exploring the Number of Trainable Blocks
To assess the impact of varying the number of trainable blocks in the ViT architecture, we conduct experiments with the last 4, 6 (as detailed in the main text), and 8 transformer blocks set as trainable. Due to computational constraints, we focus on the first three stages of MCL. All other settings remained consistent with those described in the main text. As summarized in Tab. 10, configuring 4 blocks as trainable results in an average score of 28.1, surpassing the 6-block configuration, which achieves a 27.4 average score. However, extending to 8 trainable blocks decreases the average score to 25.9. Despite these variations, MCL consistently improves performance compared to the vanilla OpenAI pretrained ViT CLIP, which scores an average of 20.0.
Table 10: Performance comparison of MCL-tuned ViT CLIP models with different numbers of trainable blocks on the MMVP benchmark. “S0” denotes the original OpenAI pre-trained ViT CLIP model. “S1” and “S2” denote Stage 1 and Stage 2. “MCL” represents the results derived from concatenating representations across the three stages.
Trainable Blocks O&D PSF S&C Q&C P&R C&A S&P Texts V&P Average
- S0 26.7 13.3 26.7 6.7 6.7 40.0 26.7 13.3 20.0 20.0
4 S1 0.0 20.0 40.0 20.0 6.7 73.3 26.7 6.7 13.3 23.0
S2 20.0 20.0 40.0 13.3 20.0 73.3 26.7 0.0 20.0 25.9
MCL 13.3 33.3 60.0 20.0 13.3 73.3 13.3 6.7 20.0 28.1
6 S1 13.3 33.3 66.7 26.7 13.3 53.3 20.0 13.3 26.7 29.6
S2 0.0 13.3 46.7 40.0 6.7 53.3 20.0 13.3 20.0 23.7
MCL 6.7 13.3 53.3 26.7 13.3 66.7 40.0 0.0 26.7 27.4
8 S1 13.3 20.0 46.7 6.7 13.3 66.7 20.0 20.0 20.0 25.2
S2 6.7 33.3 40.0 0.0 13.3 53.3 6.7 13.3 20.0 20.7
MCL 13.3 33.3 33.3 13.3 6.7 60.0 26.7 20.0 26.7 25.9
0.A.7 Clustering Analysis
Table 11: AMI scores between K-means clustering outcomes across different MCL stages. Low AMI scores between distinct stages suggest substantially different clustering results.
AMI Stage 0 Stage 1 Stage 2
Stage 0 1.00 0.33 0.25
Stage 1 0.33 1.00 0.24
Stage 2 0.25 0.24 1.00
In this section, we explore the clustering dynamics observed during the training of ViT CLIP models using the MCL approach, as detailed in Section 5.4 of the main text. We begin by examining the Adjusted Mutual Information (AMI) [49] scores computed between the clustering outcomes of subsequent stages in the experiment. AMI scores range from 0, indicating no mutual information (independent clusterings), to 1, denoting identical clustering results.
As illustrated in Tab. 11, the low AMI scores between distinct stages highlight the divergent clustering outcomes, underscoring that each stage learns a unique feature distribution. This divergence indicates the MCL framework’s effectiveness in guiding the model to capture distinct, non-redundant features across stages.
Additionally, the distribution of pseudo labels for each stage, depicted in Fig. 7 (log scale), exhibits a long-tail distribution. This pattern aligns with characteristic distributions observed in large-scale datasets [1, 46], highlighting the natural variety of data captured at different stages of MCL.
(a) Stage 0
(b) Stage 1
(c) Stage 2
Figure 7: Distribution of pseudo labels across stages 0, 1, and 2, demonstrating the variation in learned features.